### v2

@@ 4.5 Full Proof of Theorem 1
We have $\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k))\geq\frac{\sigma(\hat{\pi}^{*}(k))-\sigma(\pi^{\prime}(k))}{B}-(\delta+\Lambda\xi)$.
Proof of Lemma 1.
Given a round $t\in[T]$ and $i\in[n]$, let $\bm{v}_{t,i}$ denote the (random) $i$-th item selected at round $t$ by the greedy single-round policy ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ and let $\bm{S}_{t,i}$ denote the (random) set $\{\bm{v}_{t,1},\ldots,\bm{v}_{t,i}\}$. We recall that $\Delta_{t,i}=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)]$ is the expected increment of the objective function at round $t$ achieved by ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ when selecting the $i$-th item.
In the remainder of the proof, we will show several inequalities (Parts i-iv) that will lead to the claim (Final part).
Part i:
Given a round $t\in[T]$ and $i\in[n]$, let $\bm{v}_{t,i}^{*}$ be the (random) item that maximizes $\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)$ and $\Delta^{*}_{t,i}:=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)]$. Furthermore, let $A_{t,i}$ denote the event “$\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)\leq\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta$”; by the property (i) of Proposition 2, we have that event $A_{t,i}$ is true with probability at least $1-\xi$, and then, its complementary event $\overline{A_{t,i}}$ is false with probability at most $\xi$. Thus, we have
$\Delta^{*}_{t,i}$ $=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)]$ (13)
$=\mathbb{P}[A_{t,i}]\cdot\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)\ |\ A_{t,i}]+\mathbb{P}\left[\overline{A_{t,i}}\right]\cdot\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}\left[\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)\ |\ \overline{A_{t,i}}\right]$
$\leq\mathbb{P}[A_{t,i}]\cdot\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta\ |\ A_{t,i}]+\mathbb{P}\left[\overline{A_{t,i}}\right]\cdot\Lambda$
$\leq\mathbb{P}[A_{t,i}]\cdot\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta\ |\ A_{t,i}]+\xi\Lambda$
$\leq\mathbb{P}[A_{t,i}]\cdot\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta\ |\ A_{t,i}]+\mathbb{P}[\overline{A}_{t,i}]\cdot\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta\ |\ \overline{A}_{t,i}]+\xi\Lambda$
$=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta]+\xi\Lambda$
$=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)]+\delta+\xi\Lambda$
$=\Delta_{t,i}+\delta+\xi\Lambda.$
Part ii:
Now, let $t$ be the round to which the $(k+1)$-th unit of budget is assigned by ${\sf BudgetGr}^{\prime}_{\delta,\xi}(I)$; then, $\vec{b}^{\prime}(k)=(b_{1}^{\prime}(k),\ldots,b_{T}^{\prime}(k))$ is the budget vector obtained before the above assigned. For any round $q\neq t$, we have that
$\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k))=\Delta_{t,b^{\prime}_{t}(k)+1}$ $\geq\Delta_{q,b^{\prime}_{q}(k)+1}\geq\Delta^{*}_{q,b^{\prime}_{q}(k)+1}-(\delta+\xi\Lambda),$ (14)
where the first inequality holds by the greedy choice of ${\sf BudgetGR}^{\prime}_{\delta,\xi}(I)$ and the second one follows from (13).
Part iii:
For any $h\in[\min\{B,Tn-k\}]$, let $\hat{\Delta}^{*}_{h}(k)$ denote the expected increment of the $h$-th item selected by $\hat{\pi}^{*}(k)$ during the second phase. By construction, we have that $\sum_{h\in[\min\{B,Tn-k\}]}\hat{\Delta}^{*}_{h}(k)$ is equal to the expected increment achieved in the second phase of $\hat{\pi}^{*}(k)$, that is,
$\sum_{h\in[\min\{B,Tn-k\}]}\hat{\Delta}^{*}_{h}(k)=\sigma(\hat{\pi}^{*}(k))-\sigma(\pi^{\prime}(k))).$ (15)
Part iv:
For a given $h\in[\min\{B,Tn-k\}]$, let $\hat{\bm{t}}_{h}$ be the (random) round in which the $h$-th item of the second phase is selected by $\hat{\pi}^{*}(k)$. By exploiting the adaptive submodularity, we can show that $\Delta^{*}_{\hat{\bm{t}}_{h},b^{\prime}_{\hat{\bm{t}}_{h}}(k)+1}$ (i.e., the expected increment of the objective functions obtained by selecting optimally an item in round $\hat{\bm{t}}_{h}$, at the end of the first phase of $\hat{\pi}^{*}(k)$) is at least equal as $\hat{\Delta}^{*}_{h}(k)$ (i.e., the expected increment of the objective functions achieved by $\hat{\pi}^{*}(k)$ when selecting the $h$-th item during the second phase), independently on the particular realization of $\hat{\bm{t}}_{h}$. Indeed, let $\langle\bm{S}\rangle$ denote the partial state in round $\hat{\bm{t}}_{h}$, observable immediately after the first phase, let $\bm{v}^{*}$ be an item that can be optimally selected in the same round after observing $\langle\bm{S}\rangle$, and let $\langle\hat{\bm{S}}\rangle$ denote the partial state in the same round, but observed immediately before the selection of the $h$-th item $\hat{\bm{v}}^{*}$ in the second phase of $\hat{\pi}^{*}(k)$. We have
$\Delta^{*}_{\hat{\bm{t}}_{h},b^{\prime}_{\hat{\bm{t}}_{h}}(k)+1}=\mathbb{E}_{\langle\bm{S}\rangle}[\Delta(\bm{v}^{*}|\langle\bm{S}\rangle)]\geq\mathbb{E}_{\langle\hat{\bm{S}}\rangle}[\Delta(\hat{\bm{v}}^{*}|\langle\bm{S}\rangle)]\geq\mathbb{E}_{\langle\hat{\bm{S}}\rangle}[\Delta(\hat{\bm{v}}^{*}|\langle\hat{\bm{S}}\rangle)]=\hat{\Delta}^{*}_{h}(k),$ (16)
where the first inequality holds by the optimality of $\bm{v}^{*}$ and the last one holds since $\langle\bm{S}\rangle\preceq\langle\hat{\bm{S}}\rangle$ and by the adaptive submodularity.
We conclude that
$\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k))\geq\Delta^{*}_{\hat{\bm{t}}_{h},b^{\prime}_{\hat{\bm{t}}_{h}}(k)+1}-(\delta+\xi\Lambda)\geq\hat{\Delta}^{*}_{h}(k)-(\delta+\xi\Lambda),$ (17)
where the first inequality holds by (14) (with $q:=\hat{\bm{t}}_{h}$) and the last one holds by (16).
Final Part:
By putting together (15) and (17), we obtain
$B(\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k)))$ $\geq\min\{B,Tn-k\}\cdot(\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k)))$
[…]
$\geq\sum_{h\in[\min\{B,Tn-k\}]}\hat{\Delta}^{*}_{h}(k)-B(\delta+\xi\Lambda)$
[…]
and, by dividing all members by $B$, the claim of the lemma follows. ∎ By Lemma 1, we obtain
$\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k))\geq\frac{\sigma(\hat{\pi}^{*}(k))-\sigma(\pi^{\prime}(k))}{B}-(\delta+\xi\Lambda)\geq\frac{\sigma(\pi^{*})-\sigma(\pi^{\prime}(k))}{B}-(\delta+\xi\Lambda)$
where the last inequality follows from (12). By rearranging the above inequalities we obtain
$\sigma(\pi^{\prime}(k+1))\geq\left(1-\frac{1}{B}\right)\sigma(\pi^{\prime}(k))+\left(\frac{1}{B}\right)\sigma(\pi^{*})-(\delta+\xi\Lambda).$ (18)
By applying recursively (18) from $k=0$ to $k=B-1$ we obtain $\sigma(\pi^{\prime}(k))\geq\left(\frac{1}{B}\right)\sigma(\pi^{*})\sum_{k=0}^{B-1}\left(1-\frac{1}{B}\right)^{k}-B(\delta+\xi\Lambda)$. Thus, we have
${\sf MultiGr}^{\prime}_{0,0}(I)$ $=\sigma(\pi^{\prime}(B))$
$\geq\left(\frac{1}{B}\right)\sigma(\pi^{*})\sum_{k=0}^{B-1}\left(1-\frac{1}{B}\right)^{k}-B(\delta+\xi\Lambda)$
$=\left(\frac{1}{B}\right)\sigma(\pi^{*})\left(\frac{1-\left(1-\frac{1}{B}\right)^{B}}{1-\left(1-\frac{1}{B}\right)}\right)-B(\delta+\xi\Lambda)$
$=\left(1-\left(1-\frac{1}{B}\right)^{B}\right)\sigma(\pi^{*})-B(\delta+\xi\Lambda)$
$\geq\left(1-\frac{1}{e}\right)\sigma(\pi^{*})-B(\delta+\xi\Lambda),$ (19)
where (19) follows from the known inequality $(1+x)^{r}<e^{rx}$, holding for any $x\geq-1$ and $r>0$, applied with $x=-1/B$ and $r=B$. By (19), we have that ${\sf MultiGr}^{\prime}_{0,0}(I)$ is a $(1-1/e)$-approximate policy.
In the following lemma, we relate the value of ${\sf MultiGr}_{\delta,\xi}(I)$ with that of its variant ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$; this result, combined with (19), will be used to show the approximation factor of ${\sf MultiGr}_{\delta,\xi}(I)$.
@@
We have $\sigma({\sf MultiGr}_{\delta,\xi}(I))\geq\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))-3B(\delta+\xi\Lambda)$.
Proof of Lemma 2.
As in Lemma 1, we will first show several inequalities (Parts i-v), and then we will combine them to show the claim (Final Part).
Part i:
We recall that $\Delta_{t,i}$ denotes the expected increment of the objective function considered in algorithm ${\sf SingleGr}_{\delta,\xi}(I,t,n)$, when selecting the $i$-the item at round $t$; by construction of policy ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$ and budget vector $\vec{b}^{\prime}=(b_{1}^{\prime},\ldots,b^{\prime}_{T})$, we have that
$\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))=\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\Delta_{t,i}.$ (20)
Let $\vec{\bm{b}}=(\bm{b}_{1},\ldots,\bm{b}_{T})$ be the (random) vector returned by ${\sf BudgetGr}_{\delta,\xi}(I)$; by construction, we have that
$\sigma({\sf MultiGr}_{\delta,\xi}(I))=\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}{\Delta}_{t,i}\right],$ (21)
where the expectation is over the possible realizations of $\vec{\bm{b}}$.
Part ii:
By exploiting similar arguments as in Parts i and iv of Lemma 1, we can show that the $\Delta_{t,i}$ is non-increasing in $i$ up to addend $(\delta+\xi\Lambda)$, that is,
$\Delta_{t,i^{\prime}}\geq\Delta_{t,i}-(\delta+\xi\Lambda)$ (22)
for any $t\in[T]$ and $i,i^{\prime}\in[n]$ such that $i^{\prime}\leq i$. Indeed, let $\Delta_{t,i}^{*}$ be the maximum expected increment obtained by selecting an item at round $t$ in an optimal way, assuming that $i-1$ items have been already selected in that round by policy ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ (see Part i of the proof of Lemma 1 for a formal definition of $\Delta_{t,i}^{*}$). Inequality (22) can be shown by observing that $\Delta_{t,i^{\prime}}+\delta+\xi\Lambda\geq\Delta_{t,i^{\prime}}^{*}\geq\Delta_{t,i}^{*}\geq\Delta_{t,i}$, where the first inequality holds because of the approximate greedy choice of ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ (similarly as in (13)) and the second one holds by the adaptive submodularity (similarly as in (16)).
Part iii:
Now, let $\tilde{\bm{\Delta}}_{t,i}$ be the random quantity considered in algorithm Oracle2 to estimate $\Delta_{t,i}$ with high probability and arbitrarily high precision. Let $A$ be the event “$|\Delta_{t,i}-\tilde{\bm{\Delta}}_{t,i}|\leq\delta,\ \forall t\in[T],i\in[n]$”; by the property (i) of Proposition 3, we have $\mathbb{P}[A]\geq 1-\xi$. Then, given a realization $(\tilde{\Delta}_{t,i})_{t\in[T],i\in[n]}\sim(\tilde{\bm{\Delta}}_{t,i})_{t\in[T],i\in[n]}$ such that event $A$ is true we have that
$|\Delta_{t,i}-\tilde{\Delta}_{t,i}|\leq\delta.$ (23)
Part iv:
Let $\overline{\bm{\Delta}}_{t,i}:=\min_{i^{\prime}\in[i]}\tilde{\bm{\Delta}}_{t,i^{\prime}}$ be the (random) values considered in ${\sf BudgetGr}_{\delta,\xi}(I)$ to assign the budget in a greedy way. Given a realization $(\overline{\Delta}_{t,i})_{t\in[T],i\in[n]}\sim(\overline{\bm{\Delta}}_{t,i})_{t\in[T],i\in[n]}$ of the above values, and the corresponding budget vector $\vec{b}=(b_{1},\ldots,b_{T})$ returned as output by ${\sf BudgetGr}_{\delta,\xi}(I)$, we have that
$\sum_{t\in[T]}\sum_{i\in[b_{t}]}\overline{\Delta}_{t,i}\geq\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\overline{\Delta}_{t,i}.$ (24)
Indeed, as $\overline{\Delta}_{t,i}$ is non increasing in $i$ (by definition of minimum), one can easily show that the vector $\vec{b}^{*}=(b_{1}^{*},\ldots,b_{T}^{*})$ maximizing $\sum_{i\in[b_{t}^{*}]}\overline{\Delta}_{t,i}$ subject to $\sum_{t\in[T]}b_{t}^{*}=B$ can be computed in a greedy way as in algorithm ${\sf BudgetGr}_{\delta,\xi}(I)$, that is, the budget vector $\vec{b}^{*}$ can be set equal to $\vec{b}$, and then (24) holds by the optimality of $\vec{b}$.
Part v:
Let $(\tilde{\Delta}_{t,i})_{t\in[T],i\in[n]}$ be a realization of $(\tilde{\bm{\Delta}}_{t,i})_{t\in[T],i\in[n]}$ such that event $A$ is true, and let $\overline{\Delta}_{t,i}$ be the corresponding realization of $\overline{\bm{\Delta}}_{t,i}=\min_{i^{\prime}\in[i]}\tilde{\bm{\Delta}}_{t,i^{\prime}}$. Given $t\in[T],i\in[n]$, let $i^{\prime}\in[i]$ be the lowest integer such that $\tilde{\Delta}_{t,i^{\prime}}=\overline{\Delta}_{t,i}$. By exploiting (22), we get:
$\Delta_{t,i}-\overline{\Delta}_{t,i}$ $=\Delta_{t,i}-\tilde{\Delta}_{t,i^{\prime}}$
$\leq\Delta_{t,i^{\prime}}+\delta+\xi\Lambda-\tilde{\Delta}_{t,i^{\prime}}$ (25)
$\leq 2\delta+\xi\Lambda,$ (26)
where (25) holds by (22) and (26) holds by (23).
Final Part:
We recall that $A$ denotes the event “$|\Delta_{t,i}-\tilde{\bm{\Delta}}_{t,i}|\leq\delta,\ \forall t\in[T],i\in[n]$” and that $\mathbb{P}[A]\geq 1-\xi$; furthermore, let $\overline{A}$ denote the complementary set of $A$, so that $\mathbb{P}[\overline{A}]\leq\xi$. We conclude the proof by using all the above inequalities as follows:
[…]
By (27), the claim follows. ∎ By exploiting the polynomial boundedness property, the following lemma provides a lower bound on the optimal adaptive value.

### v3

@@ 4.6 Technical Lemmas
For any round $t\in[T]$ and real number $d\geq 0$, we have $\overline{OPT}_{t}(d)\leq 2\cdot OPT_{t}(d)$.
Proof Sketch.
We first assume that $d=b$ for some integer $b\geq 0$. The proof for a generic real number $d\geq 0$ is given in Appendix D. If $b=0$ the claim trivially follows, as $\overline{OPT}_{t}(0)=OPT_{t}(0)=0$. Thus, we assume that $b\geq 1$. Let $\pi^{*}_{t}(b)$ be an arbitrary optimal adaptive policy at round $t$ that selects $b$ items in expectation. For any $i\in[n]$, let ${\bm{S}_{i}^{\prime}}$ denote the (possibly empty) random set of at most $b$ items selected by $\pi^{*}_{t}(b)$ after having already selected $b\cdot(i-1)$ items (that is, $\sqcup_{i=1}^{n}\bm{S}_{i}^{\prime}$ is the set of items selected by the policy $\pi^{*}_{t}(b)$), $A_{i}$ denote the event “$|{\bm{S}_{i}^{\prime}}|>0$” and $\Delta(\pi^{*}_{t}(b)|A_{i})$ denote the expected increment of $f_{t}$ gained from the selection of ${\bm{S}_{i}^{\prime}}$ (after the set $\sqcup_{j=1}^{i-1}\bm{S}_{j}^{\prime}$ has already been selected) and conditioned by event $A_{i}$; furthermore, for $i\geq 2$, let $C_{i}$ denote the event “$|{\bm{S}_{i-1}^{\prime}}|=b$”. We have
$\overline{OPT}_{t}(b)=\sigma(\pi^{*}_{t}(b))=\sum_{i=1}^{n}\mathbb{P}[A_{i}]\cdot\Delta(\pi^{*}_{t}(b)|A_{i})\leq\sum_{i=1}^{n}\mathbb{P}[A_{i}]\cdot OPT_{t}(b),$ (11)
where: the second equality is obtained by decomposing the value of policy $\sigma(\pi^{*}_{t}(b))$ into a sum of expected increments and by observing that $\bm{S}_{i}^{\prime}$ is non-empty only if event $A_{i}$ is true; the last inequality holds by the strong adaptive submodularity. Furthermore, we have that [^4] We use #$\#X$ in place of $|X|$ to denote the cardinality of a set $X$. This is done to avoid confusion when symbol |||| is used for conditional expectation.
$b$ $\geq\sum_{i=2}^{n}\mathbb{E}_{\bm{\eta}_{t}}[\#{\bm{S}_{i-1}^{\prime}}]$
$\geq\sum_{i=2}^{n}\mathbb{P}[C_{i}]\cdot\mathbb{E}_{\bm{\eta}_{t}}[\#{\bm{S}_{i-1}^{\prime}}\ |\ C_{i}]$
$=\sum_{i=2}^{n}\mathbb{P}[C_{i}]\cdot b$ (12)
$\geq\sum_{i=2}^{n}\mathbb{P}[A_{i}]\cdot b$ (13)
$\geq-b+\sum_{i=1}^{n}\mathbb{P}[A_{i}]\cdot b,$ (14)
where: (12) holds since, by definition of event $C_{i}$, the cardinality of ${\bm{S}_{i-1}^{\prime}}$ is constantly equal to $b$ if event $C_{i}$ is true; (13) holds since $\mathbb{P}[C_{i}]\geq\mathbb{P}[A_{i}]$ (as event $A_{i}$ implies event $C_{i}$). By rearranging (14), we obtain
$\sum_{i=1}^{n}\mathbb{P}[A_{i}]\leq 2,$ (15)
and by applying this inequality to (11) we obtain
$\overline{OPT}_{t}(b)\leq\sum_{i=1}^{n}\mathbb{P}[A_{i}]\cdot OPT_{t}(b)\leq 2\cdot OPT_{t}(b),$ (16)
that shows the claim. ∎
@@
For any vector $\vec{d}=(d_{1},\ldots,d_{T})$ of non-negative real numbers such that $\sum_{t\in[T]}d_{t}=B$, there exists a budget vector $\vec{b^{*}}=(b_{1}^{*},\ldots,b_{T}^{*})$ of non-negative integers with $\sum_{t\in[T]}b^{*}_{t}=B$ such that
$\sum_{t\in[T]}OPT_{t}(d_{t})\leq\sum_{t\in[T]}OPT_{t}(b^{*}_{t}).$
Proof.
Let $\vec{d}=(d_{1},\ldots,d_{T})$ be a vector of non-negative real numbers such that $\sum_{t\in[T]}d_{t}=B$. For any $t\in[T]$, we have that $OPT_{t}(d_{t})=p_{t}\cdot OPT_{t}(\lceil d_{t}\rceil)+(1-p_{t})\cdot OPT_{t}(\lfloor d_{t}\rfloor)$, for $p_{t}=(d_{t}-\lfloor d_{t}\rfloor)$. We say that a round $t$ is integral if $d_{t}$ is integer, while it is fractional if $d_{t}$ is not. We observe that $p_{t}=0$ if $t$ is integral and $p_{t}\in(0,1)$ otherwise.
If all rounds are integral, we directly can set $\vec{b^{*}}=\vec{d^{*}}$ to obtain the claim. Henceforth, we assume that there exists at least a fractional round $t_{1}\in[T]$. We claim that there must necessarily be another fractional round $t_{2}\neq t_{1}$. Indeed, if all rounds $t\neq t_{1}$ were integral we would have that each $d_{t}$, with $t\neq t_{1}$, would be integer. As $\sum_{t\in[T]}d_{t}=B$, we have that $d_{t_{1}}=B-\sum_{t\neq t_{1}}d_{t}$. Thus, $d_{t_{1}}$ must be an integer, as it is defined by sums and differences of integer numbers. But this contradicts the assumption that $t_{1}$ is fractional. We conclude that that there exist at least two fractional rounds $t_{1},t_{2}$.
Assume w.l.o.g. that $10<d_{t_{1}}-\lfloor d_{t_{1}}\rfloor\leq\lceil d_{t_{2}}\rceil-d_{t_{2}}<$ 1 (otherwise, it is sufficient to permute indexes $t_{1}$ and $t_{2}$). For any $\beta\in[-(\lceil d_{t_{2}}\rceil-d_{t_{2}}),d_{t_{1}}-\lfloor d_{t_{1}}\rfloor]$, let $\vec{d}(\beta)$ be the vector obtained by moving an amount $\beta$ (resp. $-\beta$) from $d_{t_{1}}$ to $d_{t_{2}}$ if $\beta\geq 0$ (resp. from $d_{t_{2}}$ to $d_{t_{1}}$ if $\beta<0$), i.e., $d_{t_{1}}(\beta)=d_{t_{1}}-\beta$, $d_{t_{2}}(\beta)=d_{t_{2}}+\beta$ and $d_{t}(\beta)=d_{t}$ for any $t\notin\{t_{1},t_{2}\}$. Let $\Delta_{\beta}:=\beta(OPT_{t_{2}}(\lceil d_{t_{2}}\rceil)-OPT_{t_{2}}(\lfloor d_{t_{2}}\rfloor)-OPT_{t_{1}}(\lceil d_{t_{1}}\rceil)+OPT_{t_{1}}(\lfloor d_{t_{1}}\rfloor))$. For any $\beta\in[-(\lceil d_{t_{2}}\rceil-d_{t_{2}}),d_{t_{1}}-\lfloor d_{t_{1}}\rfloor]$, we have
$OPT(\vec{d}(\beta))$
$=(d_{t_{2}}+\beta-\lfloor d_{t_{2}}+\beta\rfloor)OPT_{t_{2}}(\lceil d_{t_{2}}+\beta\rceil)+(\lceil d_{t_{2}}+\beta\rceil-(d_{t_{2}}+\beta))OPT_{t_{2}}(\lfloor d_{t_{2}}+\beta\rfloor)$
$\quad+(d_{t_{1}}-\beta-\lfloor d_{t_{1}}-\beta\rfloor)OPT_{t_{1}}(\lceil d_{t_{1}}-\beta\rceil)+(\lceil d_{t_{1}}-\beta\rceil-(d_{t_{1}}-\beta))OPT_{t_{1}}(\lfloor d_{t_{1}}-\beta\rfloor)$
$\quad+\sum_{t\notin\{t_{1},t_{2}\}}OPT_{t}(d_{t})$
$=(d_{t_{2}}+\beta-\lfloor d_{t_{2}}\rfloor)OPT_{t_{2}}(\lceil d_{t_{2}}\rceil)+(\lceil d_{t_{2}}\rceil-(d_{t_{2}}+\beta))OPT_{t_{2}}(\lfloor d_{t_{2}}\rfloor)$
$\quad+(d_{t_{1}}-\beta-\lfloor d_{t_{1}}\rfloor)OPT_{t_{1}}(\lceil d_{t_{1}}\rceil)+(\lceil d_{t_{1}}\rceil-(d_{t_{1}}-\beta))OPT_{t_{1}}(\lfloor d_{t_{1}}\rfloor)$
$\quad+\sum_{t\notin\{t_{1},t_{2}\}}OPT_{t}(d_{t})$ (17)
$=\beta(OPT_{t_{2}}(\lceil d_{t_{2}}\rceil)-OPT_{t_{2}}(\lfloor d_{t_{2}}\rfloor)-OPT_{t_{1}}(\lceil d_{t_{1}}\rceil)+OPT_{t_{1}}(\lfloor d_{t_{1}}\rfloor))$
$\quad+(d_{t_{2}}-\lfloor d_{t_{2}}\rfloor)OPT_{t_{2}}(\lceil d_{t_{2}}\rceil)+(\lceil d_{t_{2}}\rceil-d_{t_{2}})OPT_{t_{2}}(\lfloor d_{t_{2}}\rfloor)$
$\quad+(d_{t_{1}}-\lfloor d_{t_{1}}\rfloor)OPT_{t_{1}}(\lceil d_{t_{1}}\rceil)+(\lceil d_{t_{1}}\rceil-d_{t_{1}})OPT_{t_{1}}(\lfloor d_{t_{1}}\rfloor)$
$\quad+\sum_{t\notin\{t_{1},t_{2}\}}OPT_{t}(d_{t})$
$=\Delta_{\beta}+OPT(\vec{d}),$ (18)
where (17) holds since, by rearranging the assumptions $\beta\in[-(\lceil d_{t_{2}}\rceil-d_{t_{2}}),d_{t_{1}}-\lfloor d_{t_{1}}\rfloor]$ and $10<d_{t_{1}}-\lfloor d_{t_{1}}\rfloor\leq\lceil d_{t_{2}}\rceil-d_{t_{2}}<$ 1, we have $\lfloor d_{t_{2}}\rfloor\leq d_{t_{2}}+\beta\leq\lceil d_{t_{2}}\rceil$ and $\lfloor d_{t_{1}}\rfloor\leq d_{t_{1}}-\beta\leq\lceil d_{t_{1}}\rceil$; furthermore, we trivially have $d_{t}(\beta)\geq 0$ for any $t\in[T]$ and $\sum_{t\in[T]}d_{t}(\beta)=B$. Finally, since $\Delta_{\beta}=0$ for $\beta=0$, we have that the maximum value of $\Delta_{\beta}$ over $\beta\in[-(\lceil d_{t_{2}}\rceil-d_{t_{2}}),d_{t_{1}}-\lfloor d_{t_{1}}\rfloor]$ is non negative; furthermore, since $\Delta_{\beta}$ is linear in $\beta\in[-(\lceil d_{t_{2}}\rceil-d_{t_{2}}),d_{t_{1}}-\lfloor d_{t_{1}}\rfloor]$, its maximum value is achieved by some $\beta\in\{-(\lceil d_{t_{2}}\rceil-d_{t_{2}}),d_{t_{1}}-\lfloor d_{t_{1}}\rfloor\}$. We conclude that there exists $\beta\in\{-(\lceil d_{t_{2}}\rceil-d_{t_{2}}),d_{t_{1}}-\lfloor d_{t_{1}}\rfloor\}$ such that $\Delta_{\beta}\geq 0$, and by (18) we get
$OPT(\vec{d}(\beta))=\Delta_{\beta}+OPT(\vec{d})\geq OPT(\vec{d});$ (19)
furthermore, as $\beta\in\{-(\lceil d_{t_{2}}\rceil-d_{t_{2}}),d_{t_{1}}-\lfloor d_{t_{1}}\rfloor\}$, we have that at least one round $t\in\{t_{1},t_{2}\}$ becomes integral in $\vec{d}(\beta)$, and the rounds that were previously integral in $\vec{d}$ continue to be integral in $\vec{d}(\beta)$. Then, the number of integral rounds of $\vec{d}(\beta)$ is higher than that related to $\vec{d}$.
We conclude that, by iteratively applying the above process to $\vec{d}^{\prime}:=\vec{d}(\beta)$, after at most $T$ iterations we obtain a vector $\vec{b}^{*}$ satisfying the claim. ∎